OpenStudy (anonymous):

i am not able to understand diffrential calculus of btech of 1 semester .please help

OpenStudy (anonymous):

hey indian? which college? me to in first year

OpenStudy (anonymous):

always try to understand how to count number of boxs in a chess board if we add one more coloumn on chess board. we generally calculate it as on a chess board we have 8 rows and 8 coloumn so the no of boxes are 8*8 = 64 if we add one more coloumn than we have to fix number of row as 8 and vary number of coloumn by 8+1=9 if we define non variable value number of row as dx (where dx = 8) and variable number of column by y. where y is variable do not have a fix value it may vary than total number of boxes = y*dx now take any value of y and calculate total no of boxes If you understand all the above discussion than you can give me an example i will elaborate it in the same manner. waiting for your reply.